### v1

# Logit Scaling for Out-of-Distribution Detection

Abstract
The safe deployment of machine learning and AI models in open-world settings hinges critically on the ability to detect out-of-distribution (OOD) data accurately, data samples that contrast vastly from what the model was trained with. Current approaches to OOD detection often require further training the model, and/or statistics about the training data which may no longer be accessible. Additionally, many existing OOD detection methods struggle to maintain performance when transferred across different architectures. Our research tackles these issues by proposing a simple, post-hoc method that does not require access to the training data distribution, keeps a trained network intact, and holds strong performance across a variety of architectures. Our method, Logit Scaling (LTS), as the name suggests, simply scales the logits in a manner that effectively distinguishes between in-distribution (ID) and OOD samples. We tested our method on benchmarks across various scales, including CIFAR-10, CIFAR-100, ImageNet and OpenOOD. The experiments cover 3 ID and 14 OOD datasets, as well as 9 model architectures. Overall, we demonstrate state-of-the-art performance, robustness and adaptability across different architectures, paving the way towards a universally applicable solution for advanced OOD detection.
keywords:
out-of-distribution detection, distribution shift, robustness
1 Introduction
Figure 1: Overview of the LTS method. LTS works at inference time during forward pass. It takes features representations and computes sample-specific scalar value which is then used to scale the logits. Final OOD detection score is calculated by applying scoring function to scaled logits. LTS incurs minimal computational costs and it doesn’t modify activations in any way thus completely preserves original working of the network while enhancing OOD detection significantly.
In the forthcoming era of artificial intelligence (AI), ensuring the safety and reliability of AI models is of utmost importance. One of the key components in achieving this is the effective detection of out-of-distribution (OOD) data on which the trained models usually severely underperform [1, 2]. Such underperformance poses a significant challenge to model reliability and safety. This is particularly crucial in areas such as healthcare, autonomous driving, and finance where the ability to detect OOD instances can profoundly influence outcomes. Thus, ensuring models can accurately recognize and manage OOD samples not only enhances their performance but also plays a vital role in preventing potential errors in critical applications.
The flip side of OOD is in-distribution (ID) data, which usually correspond to the exact training data of the model, or data samples from the same distribution as the training data. AI and machine learning research preceding the current phase of large models rely on standardized training sets that are made open and accessible, but the situation has changed starting from the large models era, where the majority of state-of-the-art models have their training data undisclosed or difficult to access. In light of this change, we are in greater need of OOD detection methods that do not necessarily require full access to the model’s training data.
The line of work in OOD detection has advanced drastically in recent years. ReAct [3] first revealed that the activation patterns in the penultimate layer of networks conceal information valuable for identifying out-of-distribution samples. ReAct [3] achieved strong OOD performance by clipping activations of the penultimate layer utilizing training data statistics. Furthermore, the methods ASH [4] and SCALE [5] demonstrated that adjusting activations in the penultimate layer of the network by pruning and scaling on a per-sample basis leads to strong OOD detection performance without the access to in-distribution data. Both methods delivered robust OOD performance, yet altering activations also resulted in a slight reduction in model accuracy. Moreover, neither of said methods demonstrated successful OOD detection over a diverse set of neural network architectures. Another recent line of work, ATS [6], has demonstrated that applying a per-sample scaling factor, calculated from training data statistics, to scale logits presents an effective method for out-of-distribution detection when applied as an addition to some of the well estabilished methods such as ReAct, ASH, Dice etc.
Method Works without training data access Doesn’t require OOD preparation steps Doesn’t affect ID accuracy
REACT [3]
DICE [7]
ASH [4]
SCALE [5]
OptFS [8]
ATS [6]
LTS (Ours)
Table 1: Preferable properties of a OOD detection methods. Table compares key properties of various OOD detection methods, highlighting whether each method operates without training data access, avoids the need for OOD-specific preparation steps (which usually involve deriving statistics from training data), and maintains in-distribution accuracy. Our proposed method, LTS, satisfies all three criteria.
In this paper, we present a logit scaling (LTS [^1] Code is available at: https://github.com/andrijazz/lts) method for out-of-distribution detection. Our method operates completely post-hoc, meaning that model does not rely on the statistics derived from the training data. This method is ready to use off-the-shelf, offering a remarkably simple, flexible and cost-effective solution. LTS leverages insights from Sun et al. [3], Djurisic et al. [4], Xu et al. [5] and relies on the feature representation from the penultimate layer. Based on that respresentation, it computes a scaling factor on a per-sample basis by using the relationship between strong and week activations. Similarly to ATS [6] the scaling factor is used to scale the logits. Once scaled, logits are then used to compute the final OOD score using well established OOD scoring function, energy score [9]. Figure 1 summarizes the inner workings of our method.
Numerous previous methods for OOD detection were constrained to a limited set of architectures. Recent research by Zhao et al. [8] conducted a thorough analysis of this issue, revealing that while many methods offer robust OOD performance for some architectures, they lack applicability across different architectures. Their method, OptFS [8], demonstrates remarkable results on architectures for which most previous methods struggled. We evaluated LTS across 9 distinct architectures and achieved superior performance compared to previous state-of-the art method, OptFS [8].
Table 1 presents an overview of the desirable properties of OOD detection methods and compares all the OOD detection methods discussed.
Our work makes the following contributions to the field of OOD detection:
  • We propose a simple and post-hoc method LTS for OOD detection that can be used off-the-shelf without access to in-distribution data.
  • We demonstrate that the LTS achieves new state-of-the-art results over 3 in-distribution (ID) and 14 out-of-distribution (OOD) datasets.
  • We demonstrate that LTS is showing robust OOD detection performance across nine different architectures, significantly reducing FPR@95 while maintaining AUROC compared to previous state-of-the-art method.
2 Related work
The field of out-of-distribution detection has significantly expanded in recent years, driven by the necessity for safer ML model deployment. Baseline methods such as those proposed by Hendrycks and Gimpel [2], Liu et al. [9], and Lee et al. [10] have laid the groundwork for much of the subsequent research in this area. We discuss four major groups of OOD detection methods relevant for understanding our own work.
OOD detection methods leveraging training data statistics. A number of earlier techniques, including ReAct [3], DICE [7], and KNN [11] employ statistics derived from training data in their OOD detection process. Specifically, ReAct adjusts the penultimate layer activations using a clipping threshold calculated from the training data. However, these methods often face challenges such as the unavailability of training data and the necessity for additional post-training processing to extract these statistics. In contrast, the proposed LTS method operates independently of training data statistics and is entirely post-hoc.
Post-hoc per sample treatment for OOD detection. In recent years, as models have become increasingly larger and their training datasets, too, the requirement to access the training data has become impractical, paving the way for methods that do not rely on the training data. ASH [4] found that treatment of activations on per-sample basis enhances OOD detection performance. Instead of implementing a global enhancement based on in-distribution (ID) data, ASH offers specific adjustments tailored to each individual sample. Follow-up work SCALE [5] also adopted this strategy, focusing on the per-sample treatment of penultimate layer activations and improved the activation treatment process. LTS employs a similar approach.
Temperature scaling. Temperature scaling, as introduced by Guo et al. [12], adjusts the logits of a neural network using a scalar value learned from a distinct validation dataset. More recently, Krumpl et al. [6] introduced the ATS, which applies temperature scaling on a per-sample basis. ATS leverages training data statistics, utilizing the activations of each sample to determine an individual scaling factor. While LTS adopts a similar approach, there are two key distinctions: firstly, unlike ATS, our method does not rely on training data statistics; secondly, LTS is an independent detection method, while ATS is used to enhance existing methods such as ReAct, ASH, and DICE.
Training time OOD enhancement. Recently, numerous approaches have been incorporating adjustments during the training phase to enhance OOD detection during inference [5, 13, 14]. These approaches typically involve either architectural modifications [13] or data augmentations during training [14]. A recent method, ISH [5], modifies the training process by introducing the scaling of activations during the backward pass. A major disadvantage of enhancements made during the training phase is the increased computational cost.
3 Logit Scaling for OOD Detection
Our method is inspired by findings reported in ReAct [3], which highlight that out-of-distribution samples often induce abnormally high activations. These findings are illustrated by Figure 2. Several methods tried to exploit this discrepancy of patterns observed in out-of-distribution and in-distribution samples [3, 4, 7, 5]. Our subsequent research, notably ASH-S [4], first observed that treating activations on a sample-based approach enhances the efficacy of OOD detectors and also lifts the requirement of accessing the training set at detection time. Still, it relied on activation pruning which resulted in modest losses of accuracy. Now we make a further step forward by proposing a method which avoids any network modification and relies only on the modification of the popular energy score.
Figure 2: Activation values and examples of ID and OOD samples. On the left we plot the activation values of all the 2048 units in the penultimate layer of a ResNet-50 pretrained on ImageNet-1k, of ID (ImageNet-1k) and OOD (iNaturalist) samples. 100 samples are taken from each dataset and the values are their average. The figure is a replication of Figure 1(b) of Sun et al. [3]. On the right we show example pictures from the corresponding dataset, including class prediction and confidence.
The energy score, as introduced by Liu et al. [9], is one of the most frequently employed techniques for detecting out-of-distribution samples. The process of detecting OOD samples using energy score operates in the following manner: raw inputs are processed by the network, the logits are computed, and fed into the energy score function. This function then produces a score used to classify whether a sample is in-distribution (ID) or out-of-distribution (OOD). The energy score is defined as a scalar value computed by the formula:
$E(\mathbf{x};f)=-\log\sum_{i=1}^{C}e^{f_{i}(\mathbf{x})}$
where $\mathbf{x}$ is the given input, $C$ is the number of classes and $f_{i}(\mathbf{x})$ is the logit for the class $i$.
Our approach modifies the aforementioned process by extending it with the computation of a scaling factor $S(\mathbf{x})$ derived from the feature representation of an individual sample $\mathbf{x}$:
$E(\mathbf{x};f)=-\log\sum_{i=1}^{C}e^{S(\mathbf{x})f_{i}(\mathbf{x})}$
The scaling factor $S$ is computed in the following manner. We denote the feature representation of an input sample $\mathbf{x}$ from the network’s penultimate layer as $h(\mathbf{x})\in\mathbb{R}^{m}$. Our method calculates the scaling factor $S$ by dividing the sum of all elements in $h(\mathbf{x})$ by the sum of the top $p$-th percent of $h(\mathbf{x})$ activations. Algorithm 1 outlines the internal mechanics of the LTS method. In the algorithm $f_{p}(\cdot)$ denotes the network up to and including the penultimate layer and $\text{logits}(\cdot)$ denotes the final layer, so the full network $f(\cdot)$ can be represented as a composition $\text{logits}\circ f_{p}$, meaning $f(\mathbf{x})=\text{logits}(f_{p}(\mathbf{x}))$ for all inputs $\mathbf{x}$. The function $\text{top}_{k}(h,k)$ returns the largest $k$ elements of $h$.
Algorithm 1 Logit Scaling for OOD detection (LTS)
Input: Single input sample $\mathbf{x}$, $\mathbf{p}$ fraction of top activations we are considering
Output: Scaling factor S
1:$h\leftarrow f_{p}(\mathbf{x})$
2:$n\leftarrow\text{dim}(h)$
3:$f\leftarrow\text{logits}(h)$
4:$S1\leftarrow\sum_{i=1}^{n}h_{i}$
5:$h_{top}=\text{top}_{k}(h,p\cdot n)$
6:$S2\leftarrow\sum_{t\in h_{top}}t$
7:$S\leftarrow\left(\frac{S1}{S2}\right)^{2}$
8:return $S$
Such simple treatment increases the scaling factor $S$ for ID samples and reduces it for OOD samples, effectively influencing separation of two distributions. The sample-based scaling factor $S$ is then used to scale the logits, facilitating the use of the energy score for the final OOD detection as shown in  Figure 1. Figure 3 illustrates the effect of applying LTS on the logits distribution and OOD detection score.
The proposed method has several desirable properties. After the forward pass is performed (which is necessary for general inference), it incurs minimal computational cost – at its core is a simple computation over the penultimate layer. Also, it is memory-efficient, requiring just a few auxiliary variables. It does not modify the activations of the network, thus preserving the network’s performance without any negative impact. It is universaly applicable, namely, it can be applied to any architecture with minimal programming effort. It performs detection based strictly on the analyzed sample and does not require the access to the data used to train the network.
Figure 3: Effect of LTS Treatment. Plots demonstrate the changes in the distribution of logits and Energy scores resulting from LTS treatment. The left-hand plots represent the state before LTS application, while the right-hand plots represent the state after LTS is applied. The plots were generated using a ResNet-50 architecture pretrained on the ImageNet-1k (ID) dataset, with iNaturalist serving as the OOD dataset. The application of LTS produces more extreme logit distribution for OOD samples and improves the separation between ID and OOD scores, leading to a substantial enhancement in OOD detection performance. The logit distribution plots were generated using a single ID and a single OOD sample, whereas the energy score plots were created using 200 images sampled from both the ID and OOD datasets.
4 Experiments
In this section, we introduce benchmarks, outline evaluation metrics, and present the experimental results of our approach, with a focus on evaluating the effectiveness of LTS and its applicability across different architectures.
4.1 OOD detection benchmark
We tested our method on four different benchmarks presented in Table 2.
CIFAR-10 benchmark ID Dataset: CIFAR-10
OOD Datasets: SVHN, LSUN C, LSUN R, iSUN, Places365, Textures
Model architecture: DenseNet-101
CIFAR-100 benchmark ID Dataset: CIFAR-100
OOD Datasets: SVHN, LSUN C, LSUN R, iSUN, Places365, Textures
Model architecture: DenseNet-101
OpenOOD benchmark ID Dataset: ImageNet-1k
OOD Datasets: SSB-hard, NINCO (Near OOD), iNaturalist, Textures, OpenImage-O (Far OOD)
Model architecture: ResNet50
ImageNet benchmark ID Dataset: ImageNet-1k
OOD Datasets: iNaturalist, SUN, Places365, Textures
Model architectures: ResNet50, MobileNetV2, ViT-B-16, ViT-L-16, SWIN-S, SWIN-B, MLP-B, MLP-L
Table 2: Benchmarks used in our OOD experiments. We evaluate our method on 4 different OOD benchmarks covering small and large scale datasets. Our tests include evaluations on 3 ID dataset, 14 OOD datasets and 9 architectures.
CIFAR-10 and CIFAR-100 benchmarks. The setup for CIFAR-10 and CIFAR-100 benchmarks is derived from Djurisic et al. [4], Sun et al. [3], Sun and Li [7] and includes 6 OOD datasets: SVHN [15], LSUN-Crop [16], LSUN-Resize [16], iSUN [17], Places365 [18] and Textures [19]. We used DenseNet-101 [20] pre-trained on corresponding ID dataset.
OpenOOD benchmark. This benchmark, introduced by  Zhang et al. [21], is designed to rigorously evaluate the effectiveness of out-of-distribution detection algorithms by providing a diverse set of datasets. OpenOOD suite divides the OOD datasets into two categories: Near-OOD and Far-OOD. The Near-OOD datasets comprise SSB-hard [22] and NINCO [23] whereas Far-OOD dataset include iNaturalist [24], Textures [19] and OpenImage-O [25]. Experiments for these tasks were conducted in alignment with setup from  Xu et al. [5] and we used ResNet50 [26] architecture pre-trained on ImageNet-1k [27].
ImageNet benchmark. The tests utilize setup derived from Djurisic et al. [4], Xu et al. [5], Zhao et al. [8]. ID dataset is ImageNet-1k and OOD datasets include iNaturalist [24], SUN [28], Places365 [18] and Textures [19]. Recent work by Djurisic et al. [4] and Xu et al. [5], evaluated ImageNet benchmark on 2 architectures ResNet50 [26], MobileNetV2 [29] while more recent study by Zhao et al. [8] extends the study to 8 different architectures. We adopted the same set of architectures and performed ImageNet benchmark evaluation on: ResNet50 [26], MobileNetV2 [29], ViT-B-16, ViT-L-16 [30], SWIN-S, SWIN-B [31], MLP-B and MLP-L [32].
4.2 OOD evaluation metrics
We evaluate our method using standard OOD detection metrics following the work of Hendrycks and Gimpel [2]: AUROC (Area Under the Receiver Operating Characteristic Curve), FPR@95 (False Positive Rate at $95\%$ True Positive Rate) and AUPR (Area Under the Precision-Recall curve). AUROC measures model’s ability to differentiate between ID and OOD classes. Higher AUROC values indicate better model performance. FPR@95 measures the proportion of false positives (incorrectly identified as positive) out of the total actual negatives, at the threshold where the true positive rate (correctly identified positives) is $95\%$. Lower FPR@95 indicates better model performance. AUPR is providing a single scalar value that reflects the model’s ability to balance precision and recall. A higher AUPR indicates a better performing model.
4.3 OOD detection performance
LTS demonstrates excellent performance in out-of-distribution detection. As highlighted in Table 3, the results on the CIFAR-10 and CIFAR-100 benchmarks show that LTS outperforms the previously leading method in both evaluation metrics. Detail CIFAR-10 and CIFAR-100 results are provided in Appendix A.
CIFAR-10 CIFAR-100
Method FPR95 AUROC FPR95 AUROC
   $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$
Softmax score 48.73 92.46 80.13 74.36
ODIN 24.57 93.71 58.14 84.49
Mahalanobis 31.42 89.15 55.37 82.73
Energy score 26.55 94.57 68.45 81.19
ReAct 26.45 94.95 62.27 84.47
DICE ${20.83}^{\pm 1.58}$ ${95.24}^{\pm 0.24}$ ${49.72}^{\pm 1.69}$ ${87.23}^{\pm 0.73}$
ASH-P 23.45 95.22 64.53 82.71
ASH-B 20.23 96.02 48.73 88.04
ASH-S 15.05 96.61 41.40 90.02
SCALE 12.57 97.27 38.99 90.74
LTS (Ours) 12.13 97.40 37.55 90.78
Table 3: OOD detection results on CIFAR benchmarks. LTS enhances the state-of-the-art performance across all evaluation metrics on the CIFAR benchmarks. The results are averaged across 6 OOD tasks. $\uparrow$ indicates that higher values are better, while a $\downarrow$ signifies that lower values are preferable. All values are presented as percentages. Table results, except for LTS (indicated as "Ours") and SCALE, are sourced from Djurisic et al. [4].
Near OOD Far OOD
Methods SSB-hard Ninco Average iNaturalist Textures OpenImage-O Average
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
$\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$
Softmax score 74.49 72.09 56.84 79.95 65.67 76.02 43.34 88.41 60.89 82.43 50.16 84.86 51.47 85.23
Mahalanobis 76.19 72.51 59.49 80.41 67.84 76.46 30.63 91.16 46.11 88.39 37.86 89.17 38.20 89.58
Energy score 76.54 72.08 60.59 79.70 68.56 75.89 31.33 90.63 45.77 88.7 38.08 89.06 38.40 89.47
ReAct 77.57 73.02 55.92 81.73 66.75 77.38 16.73 96.34 29.63 92.79 32.58 91.87 26.31 93.67
ASH-S 70.80 74.72 53.26 84.54 62.03 79.63 11.02 97.72 10.90 97.87 28.60 93.82 16.86 96.47
SCALE 67.72 77.35 51.80 85.37 59.76 81.36 9.51 98.02 11.90 97.63 28.18 93.95 16.53 96.53
LTS (Ours) 67.36 77.55 51.15 85.16 59.26 81.35 9.34 98.06 12.10 97.58 29.21 93.77 16.88 96.47
Table 4: LTS performance on OpenOOD benchmark. On the NearOOD task, LTS performs comparably to the previous state-of-the-art and slightly improves it in terms of the FPR@95 evaluation metric. On the FarOOD task, LTS performs similarly to the existing state-of-the-art method. $\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better. All values are percentages. Method results except for LTS (marked as “Ours”) are taken from Xu et al. [5].
OOD Datasets
Model Methods iNaturalist SUN Places Textures Average
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
$\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$
ResNet50 Softmax score 54.99 87.74 70.83 80.86 73.99 79.76 68.00 79.61 66.95 81.99
ODIN 47.66 89.66 60.15 84.59 67.89 81.78 50.23 85.62 56.48 85.41
Mahalanobis 97.00 52.65 98.50 42.41 98.40 41.79 55.80 85.01 87.43 55.47
Energy score 55.72 89.95 59.26 85.89 64.92 82.86 53.72 85.99 58.41 86.17
ReAct 20.38 96.22 24.20 94.20 33.85 91.58 47.30 89.80 31.43 92.95
DICE 25.63 94.49 35.15 90.83 46.49 87.48 31.72 90.30 34.75 90.77
ASH-P 44.57 92.51 52.88 88.35 61.79 85.58 42.06 89.70 50.32 89.04
ASH-B 14.21 97.32 22.08 95.10 33.45 92.31 21.17 95.50 22.73 95.06
ASH-S 11.49 97.87 27.98 94.02 39.78 90.98 11.93 97.60 22.80 95.12
ASH-B+ATS 24.07 95.19 32.70 92.37 45.63 88.33 18.71 96.29 30.28 93.05
OptFS (V) 18.33 96.63 37.03 92.84 45.97 90.15 24.80 95.48 31.53 93.78
OptFS (S) 15.90 97.00 34.00 93.28 43.61 90.50 21.61 95.99 28.78 94.19
SCALE 9.50 98.17 23.27 95.02 34.51 92.26 12.93 97.37 20.05 95.71
LTS (Ours) 9.44 98.17 22.04 95.30 32.92 92.79 15.27 96.81 19.92 95.77
MobileNet Softmax score 64.29 85.32 77.02 77.10 79.23 76.27 73.51 77.30 73.51 79.00
ODIN 55.39 87.62 54.07 85.88 57.36 84.71 49.96 85.03 54.20 85.81
Mahalanobis 62.11 81.00 47.82 86.33 52.09 83.63 92.38 33.06 63.60 71.01
Energy score 59.50 88.91 62.65 84.50 69.37 81.19 58.05 85.03 62.39 84.91
ReAct 42.40 91.53 47.69 88.16 51.56 86.64 38.42 91.53 45.02 89.47
DICE 43.09 90.83 38.69 90.46 53.11 85.81 32.80 91.30 41.92 89.60
ASH-P 54.92 90.46 58.61 86.72 66.59 83.47 48.48 88.72 57.15 87.34
ASH-B 31.46 94.28 38.45 91.61 51.80 87.56 20.92 95.07 35.66 92.13
ASH-S 39.10 91.94 43.62 90.02 58.84 84.73 13.12 97.10 38.67 90.95
SCALE 38.20 91.67 42.64 90.09 58.73 84.00 12.59 97.47 38.04 90.81
LTS (Ours) 29.78 94.33 36.14 92.34 50.65 87.84 14.02 96.89 32.65 92.85
Table 5: OOD detection results on ImageNet-1k benchmark. LTS outperforms all existing methods on this benchmark. $\uparrow$ indicates that higher values are better, while $\downarrow$ signifies that lower values are preferable. Results indicated "Ours" are computed by us. ATS results are copied from Krumpl et al. [6]. ATS performs best when combined with ASH-B, so we have included their best results for reference. The remaining values in the table are sourced from Table 1 in Djurisic et al. [4].
On the OpenOOD benchmark, LTS achieves performance comparable to the previous state-of-the-art method, SCALE [5]. More specifically, on Near-OOD benchmark LTS marginally improves FPR@95 while on Far-OOD benchmark its performance is slightly below that of the previous state-of-the-art. Table 4 showcase LTS results on OpenOOD benchmark.
Table 5 showcases the performance of LTS on the ImageNet-1k benchmark across two architectures, ResNet-50 and MobileNetV2. LTS surpasses the previous benchmarks in both metrics across all OOD tasks, with the exception of ImageNet-1k (ID) vs Textures task (OOD). On average, LTS sets a new state-of-the-art, offering improvements in both metrics for ResNet-50. For MobileNetV2, it achieves a reduction in FPR@95 by $3\%$ while enhancing the AUROC evaluation metric.
4.4 LTS applicability accross architectures
A recent study by Zhao et al. [8] revealed that many previous methods fail to maintain robust out-of-distribution performance across different architectures. Specifically, modern transformer-based architectures such as ViT [30], SwinTransformer [31], as well as MLP-Mixer [32] present significant challenges for existing OOD detection methods. The OptFS [8] method was the first to demonstrate consistent performance across eight different architectures. In our experiments, we replicate the exact experimental setup of OptFS and demonstrate that our method performs consistently well across all tested architectures, significantly reducing FPR@95 while maintaining a comparable level of AUROC compared to OptFS. Figure 4 illustrates the performance of LTS across five different architectures. Detailed results for all tested architectures can be found in Appendix B. Integration of LTS within the different architectures is described in Appendix C.
(a) AUROC evaluation (higher means better)
(b) FPR@95 evaluation (lower means better)
Figure 4: Performance comparison of OOD methods across five architectures. This figure compares the performance of various OOD detection methods across five different architectures. The evaluation is based on two metrics: (a) AUROC and (b) FPR@95. All results are tested on ImageNet benchmark and averaged across 4 tasks (iNaturalist, SUN, Places, and Textures). Higher AUROC indicates better performance, while lower FPR@95 indicates better performance.
4.5 Ablation studies
Figure 5: Estimating optimal threshold $p$ for scaling factor calculation. We evaluated LTS with various values of the hyperparameter $p$ across five architectures and averaged the results over four tasks. Our findings indicate that LTS achieves optimal performance at $p=5\%$ on both evaluation metrics, AUROC and FPR@95. Therefore, we recommend using $p=5\%$ as the default value.
Determining the optimal value of $p$. In this section we evaluate the performance of LTS across various values of $p$, which represents the proportion of top activations used in computing the scaling factor $S$. The detailed analysis is shown in Figure 5. The appropriate choice of $p$ is highly dependent upon the specific task as well as network architecture in use. In Figure 5, we present a comprehensive sweep across five architectures: ResNet-50, MobileNet-V2, ViT-16/B, Swin Transformer S and MLP-Mixer-B and 4 different tasks from ImageNet-1k benchmark. Based on our analysis, we have found that $p$ value of 5% generally yields optimal results.
Compatibility with other OOD detection methods. LTS is not compatible with methods that are relying on scaling activations in middle layers of the network due to its operational characteristics. For instance, ASH-S and SCALE perform scaling of activations on the penultimate layer, which is quite similar to the LTS operation. Consequently, combining these methods with LTS is not practical.
On the other hand, ReAct [3] greatly benefits from LTS. ReAct leverages the observation illustrated by Figure 2 to clip penultimate layer activations, thus distinguishing between in-distribution and out-of-distribution data. In our approach, we first use LTS to calculate the scaling factor $S$, then apply the ReAct rectification operation, and finally scale the logits using $S$. Table 6 presents the results of combining these two methods, showing that LTS enhances ReAct’s performance across all metrics.
Model Methods Average
FPR95 AUROC AUPR
$\downarrow$ $\uparrow$ $\uparrow$
ResNet-50 ReAct 30.72 93.27 98.59
ReAct + LTS 19.74 95.77 99.08
MobileNet ReAct 48.95 88.75 97.46
ReAct + LTS 33.57 92.59 98.29
ViT-B/16 ReAct 64.99 80.74 94.65
ReAct + LTS 57.59 84.24 96.05
Swin Transformer S ReAct 67.47 72.26 90.93
ReAct + LTS 61.97 79.99 94.58
Table 6: Compatibility of LTS and ReAct. ReAct significantly benefits when combined with LTS. Our approach involves first using LTS to calculate the scaling factor $S$, then applying the ReAct rectification operation, and finally scaling the logits using $S$. In the results table, $\uparrow$ indicates that higher values are better, while $\downarrow$ signifies that lower values are preferable. Presented results are averaged across 4 tasks on ImageNet benchmark.
5 Conslusion
In this study, we introduced LTS, extremely simple, post-hoc, off-the-shelf method for detecting out-of-distribution samples. LTS operates by deriving a scaling factor for each sample based on activations from the penultimate layer, which is then applied to adjust the logits. We conducted thorough testing of LTS, demonstrating that its performance surpasses many existing methods. Additionally, we have shown its robustness and effectiveness across a diverse set of architectures. Looking ahead, we plan to explore strategies to consistently maintain performance on specific OOD tasks irrespective of architectural differences.
Appendix A Detailed CIFAR-10 And CIFAR-100 Results
Table 7 and Table 8 supplement Table 3 in the main text, as they display the full results on each of the 6 OOD datasets for models trained on CIFAR-10 and CIFAR-100 respectively.
Method SVHN LSUN-c LSUN-r iSUN Textures Places365 Average
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
   $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$
Softmax score 47.24 93.48 33.57 95.54 42.10 94.51 42.31 94.52 64.15 88.15 63.02 88.57 48.73 92.46
ODIN 25.29 94.57 4.70 98.86 3.09 99.02 3.98 98.90 57.50 82.38 52.85 88.55 24.57 93.71
GODIN 6.68 98.32 17.58 95.09 36.56 92.09 36.44 91.75 35.18 89.24 73.06 77.18 34.25 90.61
Mahalanobis 6.42 98.31 56.55 86.96 9.14 97.09 9.78 97.25 21.51 92.15 85.14 63.15 31.42 89.15
Energy score 40.61 93.99 3.81 99.15 9.28 98.12 10.07 98.07 56.12 86.43 39.40 91.64 26.55 94.57
ReAct 41.64 93.87 5.96 98.84 11.46 97.87 12.72 97.72 43.58 92.47 43.31 91.03 26.45 94.67
DICE 25.99±5.10 95.90±1.08 0.26±0.11 99.92±0.02 3.91±0.56 99.20±0.15 4.36±0.71 99.14±0.15 41.90±4.41 88.18±1.80 48.59±1.53 89.13±0.31 20.83±1.58 95.24±0.24
ASH-P 30.14 95.29 2.82 99.34 7.97 98.33 8.46 98.29 50.85 88.29 40.46 91.76 23.45 95.22
ASH-B 17.92 96.86 2.52 99.48 8.13 98.54 8.59 98.45 35.73 92.88 48.47 89.93 20.23 96.02
ASH-S 6.51 98.65 0.90 99.73 4.96 98.92 5.17 98.90 24.34 95.09 48.45 88.34 15.05 96.61
SCALE 5.77 98.72 0.72 99.74 3.35 99.22 3.44 99.21 23.39 94.97 38.61 91.74 12.55 97.27
LTS (Ours) 5.54 98.81 0.74 99.74 3.49 99.21 3.70 99.20 21.44 95.56 37.84 91.89 12.12 97.40
Table 7: Detailed results on six common OOD benchmark datasets: Textures [19], SVHN [15], Places365 [18], LSUN-Crop [16], LSUN-Resize [16], and iSUN [17]. For each ID dataset, we use the same DenseNet pretrained on CIFAR-10. $\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better.
Method SVHN LSUN-c LSUN-r iSUN Textures Places365 Average
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
   $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$    $\downarrow$    $\uparrow$
Softmax score 81.70 75.40 60.49 85.60 85.24 69.18 85.99 70.17 84.79 71.48 82.55 74.31 80.13 74.36
ODIN 41.35 92.65 10.54 97.93 65.22 84.22 67.05 83.84 82.34 71.48 82.32 76.84 58.14 84.49
GODIN 36.74 93.51 43.15 89.55 40.31 92.61 37.41 93.05 64.26 76.72 95.33 65.97 52.87 85.24
Mahalanobis 22.44 95.67 68.90 86.30 23.07 94.20 31.38 93.21 62.39 79.39 92.66 61.39 55.37 82.73
Energy score 87.46 81.85 14.72 97.43 70.65 80.14 74.54 78.95 84.15 71.03 79.20 77.72 68.45 81.19
ReAct 83.81 81.41 25.55 94.92 60.08 87.88 65.27 86.55 77.78 78.95 82.65 74.04 62.27 84.47
DICE 54.65±4.94 88.84±0.39 0.93±0.07 99.74±0.01 49.40±1.99 91.04±1.49 48.72±1.55 90.08±1.36 65.04±0.66 76.42±0.35 79.58±2.34 77.26±1.08 49.72±1.69 87.23±0.73
ASH-P 81.86 83.86 11.60 97.89 67.56 81.67 70.90 80.81 78.24 74.09 77.03 77.94 64.53 82.71
ASH-B 53.52 90.27 4.46 99.17 48.38 91.03 47.82 91.09 53.71 84.25 84.52 72.46 48.73 88.04
ASH-S 25.02 95.76 5.52 98.94 51.33 90.12 46.67 91.30 34.02 92.35 85.86 71.62 41.40 90.02
SCALE 24.35 94.99 18.24 96.09 53.61 87.71 46.66 89.43 30.25 91.16 93.64 54.65 44.46 85.67
LTS (Ours) 21.57 96.11 4.89 99.07 43.29 91.92 38.35 93.02 31.12 93.15 85.84 71.41 37.51 90.78
Table 8: Detailed results on six common OOD benchmark datasets: Textures [19], SVHN [15], Places365 [18], LSUN-Crop [16], LSUN-Resize [16], and iSUN [17]. For each ID dataset, we use the same DenseNet pretrained on CIFAR-100. $\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better.
Appendix B LTS Performance Across Eight Architectures
Table 9 supplements Figure 4 and presents detailed performance of LTS across 8 different architectures on ImageNet benchmark, along with other methods.
[…]
caption-above
Method ResNet50 MobileNetV2 ViT-B-16 ViT-L-16 SWIN-S SWIN-B MLP-B MLP-L Average
[…]
MSP 64.76 82.82 70.47 80.67 61.74 83.12 65.22 81.75 59.68 83.75 62.79 81.38 69.36 81.97 76.01 80.04 66.25 81.94
ODIN 57.68 87.08 60.42 86.39 61.24 79.59 64.06 78.65 57.30 80.99 64.36 73.77 63.05 85.15 73.17 82.29 62.66 81.74
Energy 57.47 87.05 58.87 86.59 67.41 74.31 68.43 74.65 62.82 77.65 75.32 64.87 85.99 79.96 84.44 79.38 70.08 78.06
DICE 35.72 90.92 41.93 89.60 90.30 70.77 71.77 67.12 88.68 39.70 40.71 59.71 83.36 62.63 81.53 75.40 67.55 70.49
ReAct 30.70 93.30 50.09 88.81 67.08 81.78 59.68 83.50 50.09 87.02 64.86 83.24 90.41 77.86 80.71 79.13 63.73 84.44
BFAct 31.41 92.98 48.35 89.19 73.26 82.75 81.16 82.69 57.20 88.21 65.44 86.62 96.76 67.46 96.36 72.16 68.74 82.76
ASH-P 50.33 89.04 57.15 87.34 99.45 20.30 99.42 18.37 99.11 19.21 99.08 20.59 98.77 36.88 99.04 29.20 87.79 40.11
ASH-B 22.73 95.06 35.66 92.13 94.33 49.14 94.08 37.89 96.42 30.00 91.47 47.38 99.83 19.14 66.05 84.31 75.07 56.88
ASH-S 22.80 95.12 38.67 90.95 99.62 16.47 99.55 16.72 99.36 14.98 99.35 17.75 98.31 34.85 99.36 18.65 82.13 38.18
VRA-P 25.49 94.57 45.53 89.85 98.02 34.64 99.62 14.99 99.38 18.15 99.65 15.51 99.64 16.58 99.23 19.49 83.32 37.97
SCALE 20.05 95.71 38.04 90.81 78.34 64.08 92.77 49.25 91.95 52.21 79.07 68.75 80.62 81.35 96.87 68.91 72.21 71.38
OptFS(V) 31.53 93.78 49.09 89.62 62.01 85.80 68.61 85.10 61.47 87.04 62.96 86.09 80.23 78.83 82.59 78.92 62.31 85.65
OptFS(E) 28.78 94.19 46.92 89.90 63.04 85.64 74.96 84.69 60.67 87.50 63.50 86.55 91.37 74.85 88.49 78.22 64.71 85.19
LTS (Ours) 19.92 95.77 29.78 94.33 57.31 83.12 65.05 75.33 49.43 84.87 63.97 78.57 74.80 81.86 78.01 79.69 54.78 84.19
Table 9: OOD detection results on ImageNet benchmark across 8 model architectures. OOD datasets are iNaturalist, SUN, Places, and Textures. $\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better.
Appendix C Application of LTS to different architectures
Figure 6 illustrates the integration of LTS within the Vision Transformer and ResNet-50 architectures. We observe that ResNet-50, DenseNet-101 and MobileNetV2 have non-negative activations at penultimate layer, unlike ViT, Swin Transformer and MLP-Mixer. To mimic that, in OOD detection time we apply ReLU on penultimate layer activations of ViT, Swin Transformer and MLP-Mixer before feeding them into LTS. We emipirically observed that such a modification leads to increase in OOD detection performance.
Figure 6: Integration of LTS in Vision Transformer (ViT) and ResNet-50 architectures. The diagrams illustrate the placement of the LTS module, which calculates scaling factor used to adjust the logits before applying OOD detection scoring function. In ViT, LTS is applied before the MLP Head, following the Transformer Encoder’s output. In ResNet-50, LTS is added post average pooling layer and before the fully connected layer.